Circular No. 8115

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2003 G1 (LINEAR) An apparently asteroidal object reported by LINEAR (discovery ob-

servation below), and posted on the NEO Confirmation Page, has been found to be cometary by several CCD observers, including L. Šarounová and P. Kušnirák (Ondřejov), A. Galád (Modra), P. Birtwhistle (Great Shefford, U.K.), G. Hug (Eskridge, KS), P. R. Holvorcem (0.81-m Tenagra II telescope; m_{1} = 15.4 on Apr. 9.46 UT), and M. Tichý (Kleť). The general description of the comet gives a coma of diameter 8′′–15′′ and a straight tail ∼ 40′′–90′′ long in p.a. 210^{o}–225^{o} during Apr. 9.1–10.0. The available astrometry, preliminary parabolic orbital elements (T = 2003 Feb. 7, q = 4.9 AU, i = 67^{o}), and ephemeris appear on MPEC 2003-G56.

2003 UT  α_{2000}  δ_{2000}  m_{2}
Apr. 8.45282  17^{h}42^{m}30.^{s}50  −3^{o}33′48′′.3  17.0

SUPERNOVAE 2003dj, 2003dk, AND 2003dl Three additional apparent supernovae have been reported from unfil-

tered CCD survey images — SNe 2003dj and 2003dk by W. M. Wood-Vasey et al. (NEAT; cf. IAUC 8114), and SN 2003dl in NGC 5789 by J. Graham and W. Li (LOTOSS/KAIT; cf. IAUC 8113).

SN  2003 UT  α_{2000}  δ_{2000}  Mag.  Offset
2003dj Mar. 24.46 13^{h}16^{m}22.^{s}04 +19^{o}25′53′′.4 20.0 3′′.4 E, 2′′.1 N
2003dk Mar. 27.34 13 15 31.65 − 7 20 55.8 19.2 2′′.3 W, 3′′.1 S
2003dl Mar. 31.5 14 56 36.25 +30 14 21.5 19.2 7′′.8 E, 21′′.0 N

Additional approximate magnitudes from the respective observers: SN 2003dj, 2002 Apr. 10.27 UT, [21.7; 2003 Feb. 3.62, [20.8; Mar. 26.46, 20.2; Apr. 9.31, 19.9. SN 2003dk, 2002 May 9.16, [21.1; 2003 Feb. 8.56, [20.9; 18.61, [20.0; Apr. 9.31, 19.7. SN 2003dl, Feb. 8.5, [20.0; Mar. 16.5, [19.5; Apr. 4.4, 19.2 (M. Schwartz, Tenagra II telescope); 8.5, 19.2; 9.5, 19.2.

V4006 SAGITTARII (NO NEW NOVA IN SAGITTARIUS) The ‘probable nova’ reported on IAUC 8113 appears to be the Mira-

type variable V4006 Sgr, now near its peak brightness (period ∼ 281 days). The rough discovery position given by Lukas (1976, JAAVSO 4, 64), is ∼ 1′ southwest of the precise position given on IAUC 8113.

2003 April 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green